# bioSBM: a random graph model to integrate epigenomic data in chromatin structure prediction

### v1

@@ IV Discussion and conclusions
$f(\eta_{i})=-\frac{1}{2}(\eta_{i}-\mu_{i})^{T}(\Sigma)^{-1}(\eta_{i}-\mu_{i})+\sum_{j(\neq i)}^{N}\sum_{k=1}^{K}\left(\eta_{i,k}-\log\sum_{k^{\prime}=1}^{K}\exp(\eta_{i,k^{\prime}})\right)\mathbb{E}_{q(z_{ij})}[z_{ij,k}]$ (S6)
@@
$\nabla f(\eta_{i})=\sum_{j(\neq i)}^{N}\left(\mathbb{E}_{q(z_{ij})}[z_{ij}]-\theta(\eta_{i})\sum_{k=1}^{K}\mathbb{E}_{q(z_{ij})}[z_{ij,k}]\right)-\Sigma^{-1}(\eta_{i}-\mu_{i})$ (S7)
$\nabla^{2}f(\eta_{i})=\sum_{j(\neq i)}^{N}\left((\text{diag}(\theta(\eta_{i}))-\theta(\eta_{i})\theta(\eta_{i})^{T})\sum_{k=1}^{K}\mathbb{E}_{q(z_{ij})}[z_{ij,k}]\right)-\Sigma^{-1}$ (S8)
For the topic assignments distributions we have $q(z_{ij})=\text{Mult}(\phi_{ij})$ with $\phi_{ij}$ as variational parameters, then we can explicit the expectations in the above expressions as:
@@
Where $\text{diag}(\theta)$ is the diagonal matrix where the non-zero elements are the components of $\theta$. It can be shown that $-\nabla^{2}f(\eta_{i})$ is always a positive definite matrix, therefore the method always gives a valid covariance matrix for the variational approximation. We use the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm with second-order backtracking line-search [50]. The optimization procedure returns $(\hat{\eta}_{i},-\nabla^{2}f(\hat{\eta}_{i})^{-1})$, thereby updating the variational distribution on $\eta$, $q(\eta_{i})\simeq\mathcal{N}(\eta_{i}|\,\hat{\eta}_{i},-\nabla^{2}f(\hat{\eta}_{i})^{-1})$. This is the E-step for the membership distributions per node. This step, as for most variational mean field methods for mixed membership block models, has a cost which is $\sim\mathcal{O}(N^{2})$.
For the topic assignments posteriors we explicitly write Eq. (S3) and we get that for every $(i,j)$ and topic $k$
$\phi_{ij,k}\propto$ $\exp\left\{\sum_{g=1}^{K}\phi_{ji,g}\log P(Y_{ij}|B_{kg},\sigma_{kg})+\mathbb{E}_{q(\eta_{i})}[\log P(z_{ij}|\eta_{i})]\right\}$ (S11)
$=$ $\exp\{\mathbb{E}_{q(\eta_{i})}[\log P(z_{ij}|\eta_{i})]\}\prod_{g=1}^{K}P(Y_{ij}|B_{kg},\sigma_{kg})^{\phi_{ji,g}}$
@@
[…]
By imposing normalization $\sum_{k}\phi_{ij,k}=1$ for all pairs $(i,j)$ the updates for the topic assignments posteriors become
$\phi_{ij,k}\propto\exp(\lambda_{i,k})\prod_{g=1}^{K}P(Y_{ij}|B_{kg}\sigma_{kg})^{\phi_{ji,g}}$ (S13)
@@
$\log\phi_{ij,k}\,=\,\lambda_{i,k}-\sum_{g=1}^{K}\left[\frac{(Y_{ij}-B_{kg})^{2}}{2\sigma_{kg}}+\frac{\log\sigma_{kg}}{2}\right]\phi_{ji,g}+const.$ (S14)
Everything described until now represents the E-step of the VBEM algorithm and the updates are performed with fixed $(\Sigma,\Gamma,B,\sigma)$. The values for these parameters are updated in the M-step where instead we fix the variational parameters $\{\lambda_{i},\nu_{i}\}\text{ for }i=1\ldots N$ and $\phi_{ij}\text{ for }i\neq j$ that are optimized in the E-step.
@@
Something additional that we can do is to add a regularization to the norm of $\Sigma$ for numerical stability. In particular, something simple we can do is to regularize using the nuclear norm. For a generic matrix, the nuclear norm is given by the sum of the singular values; for a positive definite matrix such as $\Sigma$ it is easy to show that the norm corresponds to the trace of the matrix $\text{Tr}(\Sigma)$. We then introduce an additional term $-R\cdot\text{Tr}(\Sigma)$ in the ELBO where $R=rN$ (selected with cross-validation) is the regularization strength and we write the explicit $N$ dependence in such a way that all the terms of the ELBO containing $\Sigma$ are extensive. The results analyzed in the paper are obtained using $r=0.1$ With the additional term, Eq. S16 reads:
@@
=\,= $\frac{N}{2}\Sigma-\frac{1}{2}\sum_{i=1}^{N}\nu_{i}^{T}+(\lambda_{i}-\mu_{i})(\lambda_{i}-\mu_{i})^{T}+R\Sigma^{2}=0$ (S17)
@@
$\frac{\partial\mathcal{L}(q,\Psi)}{\partial B_{kg}}=\sum_{i\neq j}^{N}\phi_{ij,k}\phi_{ji,g}\frac{Y_{ij}-B_{kg}}{\sigma_{kg}}$ (S19)
@@
S2 Dataset preprocessing
Chromatin contact files were downloaded from the 4DN consortium data portal [51], and log-observed-over-expected Hi-C maps were extracted from these files using the hic-straw [52] Python toolkit. To remove anomalously low reads from the Hi-C maps, we remove all rows and columns where the average value of the normalized Hi-C matrix has a modified Z-score smaller than $-3.0$.
@@
S3 Simulation of predicted maps
@@
H3K27me3 (ChIP-seq) ENCFF717ZKL ENCSR810BDB
H3K4me1 (ChIP-seq) ENCFF239FXT ENCSR161MXP
H3K36me3 (ChIP-seq) ENCFF024LGD ENCSR091QXP
H3K9me3 (ChIP-seq) ENCFF254TIW ENCSR179BUC
HCT116 H3K9ac (ChIP-seq) ENCFF743PJP ENCSR093SHE
H3K4me2 (ChIP-seq) ENCFF693HVA ENCSR794ULT
H4K20me1 (ChIP-seq) ENCSR474DOV ENCSR474DOV
H2AFZ (ChIP-seq) ENCFF863EHT ENCSR227XNT
H3K79me2 (ChIP-seq) ENCFF631DLM ENCSR494CCN
CTCF (ChIP-seq) ENCFF787LAV ENCSR450DXU
POLR2A (ChIP-seq) ENCFF802CGI ENCSR [^0] EUU
RAD21 (ChIP-seq) ENCFF776IXR ENCSR [^0] BSB
@@
Figure S1: The Evidence Lower BOund (ELBO) was used as a criterion for Bayesian model selection. The left plot shows ELBO as a function of the number of communities of the model for the training set comprising odd-numbered chromosomes from 11 to 21. The right plot refers to even-numbered chromosomes from 12 to 22. We identify the optimal number of communities as the value where the ELBO peaks or plateaus, leading to the choice $K=12$.
Figure S2: (Left) Shannon entropies of the MAP estimates of the membership vectors for all the genomic regions under study. The two cell lines that have been studied display a similar distribution of said entropy. In both cases the peak at zero shows that a big fraction of nodes displays a single membership behavior, however, there is still a considerable degree of mixed-membershipness if one looks at the entirety of the genomic regions under study. (Right) Normalized Mutual Information (NMI) score between the clustering obtained from the model and the biological annotations. Allowing mixed membership leads to a better recapitulation of the biological annotations.
@@
Figure S3: Correlations between model-generated maps and experimental log O/E Hi-C maps for two models with a different map from biochemical covariates to membership vectors. The turquoise plots correspond to the case where $\Gamma$ is the simple $14\times 12$ matrix representing a linear mapping, discussed in the main text. The red boxes correspond to the case where $\Gamma()$ is a feed-forward deep neural network with 4 hidden layers of size $(64,64,64,32)$ with the standard ReLU (rectified linear unit) non-linear activations. While on the training set, the non-linear model better reproduces experimental contact maps, already for test chromosomes on the same cell-line the correlations decrease drastically.

### v2

@@ Acknowledgements.
$f(\eta_{i})=-\frac{1}{2}(\eta_{i}-\mu_{i})^{T}(\Sigma)^{-1}(\eta_{i}-\mu_{i})+\sum_{j(\neq i)}^{N}\sum_{k=1}^{K}\left(\eta_{i,k}-\log\sum_{k^{\prime}=1}^{K}\exp(\eta_{i,k^{\prime}})\right)q(z_{ij}=k)$ (S6)
@@
$\nabla f(\eta_{i})=\sum_{j(\neq i)}^{N}\left(\mathbb{E}_{q(z_{ij})}[z_{ij}]-\theta(\eta_{i})\sum_{k=1}^{K}q(z_{ij}=k)\right)-\Sigma^{-1}(\eta_{i}-\mu_{i})$ (S7)
$\nabla^{2}f(\eta_{i})=\sum_{j(\neq i)}^{N}\left((\text{diag}(\theta(\eta_{i}))-\theta(\eta_{i})\theta(\eta_{i})^{T})\sum_{k=1}^{K}q(z_{ij}=k)\right)-\Sigma^{-1}$ (S8)
For the community assignments distributions we have $q(z_{ij})=\text{Mult}(\phi_{ij})$ with $\phi_{ij}$ as variational parameters, then we can explicit the expectations in the above expressions as:
@@
Where $\text{diag}(\theta)$ is the diagonal matrix where the non-zero elements are the components of $\theta$. It can be shown that $\nabla^{2}f(\eta_{i})$ is always a positive definite matrix, therefore the method always gives a valid covariance matrix for the variational approximation. We use the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm with second-order backtracking line-search [66]. The optimization procedure returns $(\hat{\eta}_{i},-\nabla^{2}f(\hat{\eta}_{i})^{-1})$, thereby updating the variational distribution on $\eta$, $q(\eta_{i})\simeq\mathcal{N}(\eta_{i}|\,\hat{\eta}_{i},-\nabla^{2}f(\hat{\eta}_{i})^{-1})$. This is the E-step for the membership distributions per node. This step, as for most variational mean field methods for mixed membership block models, has a cost which is $\sim\mathcal{O}(N^{2})$.
For the community assignments posteriors we explicitly write Eq. (S3) and we get that for every $(i,j)$ and community $k$
$\phi_{ij,k}\propto$ $\exp\left\{\sum_{g=1}^{K}\phi_{ji,g}\log P(Y_{ij}|B_{kg},\sigma^{2}_{kg})+\mathbb{E}_{q(\eta_{i})}[\log P(z_{ij}|\eta_{i})]\right\}$ (S11)
$=$ $\exp\{\mathbb{E}_{q(\eta_{i})}[\log P(z_{ij}|\eta_{i})]\}\prod_{g=1}^{K}P(Y_{ij}|B_{kg},\sigma^{2}_{kg})^{\phi_{ji,g}}$
@@
$=$ $\,\,\lambda_{i,k}-\int d\eta_{i}\,\mathcal{N}(\lambda_{i},\nu_{i})\log\sum_{g=1}^{K}\exp(\eta_{i,g})$
By imposing normalization $\sum_{k}\phi_{ij,k}=1$ for all pairs $(i,j)$ the updates for the community assignments posteriors become
$\phi_{ij,k}\propto\exp(\lambda_{i,k})\prod_{g=1}^{K}P(Y_{ij}|B_{kg},\sigma^{2}_{kg})^{\phi_{ji,g}}$ (S13)
@@
$\log\phi_{ij,k}\,=\,\lambda_{i,k}-\sum_{g=1}^{K}\left[\frac{(Y_{ij}-B_{kg})^{2}}{2\sigma^{2}_{kg}}+\frac{\log\sigma^{2}_{kg}}{2}\right]\phi_{ji,g}+const.$ (S14)
Everything described until now represents the E-step of the VBEM algorithm and the updates are performed with fixed $(\Sigma,\Gamma,B,\sigma^{2})$. The values for these parameters are updated in the M-step where instead we fix the variational parameters $\{\lambda_{i},\nu_{i}\}\text{ for }i=1\ldots N$ and $\phi_{ij}\text{ for }i\neq j$ that are optimized in the E-step.
@@
Something additional that we can do is to add a regularization to the norm of $\Sigma$ for numerical stability. In particular, something simple we can do is to regularize using the nuclear norm. For a generic matrix, the nuclear norm is given by the sum of the singular values; for a positive definite matrix such as $\Sigma$ it is easy to show that the norm corresponds to the trace of the matrix $\text{Tr}(\Sigma)$. We then introduce an additional term $R\cdot\text{Tr}(\Sigma)$ in the ELBO where $R=rN$ (selected with cross-validation) is the regularization strength and we write the explicit $N$ dependence in such a way that all the terms of the ELBO containing $\Sigma$ are extensive. The results analyzed in the paper are obtained using $r=0.1$ With the additional term, Eq. (S16) reads:
@@
$=\,$ $\frac{N}{2}\Sigma-\frac{1}{2}\sum_{i=1}^{N}\nu_{i}^{T}+(\lambda_{i}-\mu_{i})(\lambda_{i}-\mu_{i})^{T}+R\Sigma^{2}=0$ (S17)
@@
$\frac{\partial\mathcal{L}(q,\Psi)}{\partial B_{kg}}=\sum_{i\neq j}^{N}\phi_{ij,k}\phi_{ji,g}\frac{Y_{ij}-B_{kg}}{\sigma^{2}_{kg}}$ (S19)
@@
S2 Method details
S2.1 Dataset preprocessing
Chromatin contact files were downloaded from the 4DN consortium data portal [67], and log-observed-over-expected Hi-C maps were extracted from these files using the hic-straw [68] Python toolkit. We have processed each arm of each chromosome separately, with the exception of the very short $p$ arms of the acrocentric chromosomes 13,14, 15, 21 and 22. To remove anomalously low reads from the Hi-C maps, we remove all rows and columns where the average value of the normalized Hi-C matrix has a modified Z-score smaller than 3.0-3.0.
@@
S2.2 Simulation of predicted maps
@@
Note that while the training of the model has been carried out looking only at intra-arm contacts data, the predicitons are made on whole chromosomes, including the contact between $p$ and $q$ arm.
S2.3 Log-fold enrichment
Given a partition of a set of $N$ elements $U=\{u_{1},\ldots,u_{N}\}$, where $u_{i}$ is the label assigned to the $i$-th element, and given a feature vector $F=\{f_{1},\ldots,f_{N}\}$, the log-fold enrichment of a certain cluster $k$ in the feature is defined as:
$\ln\Big{[}\frac{\sum_{i=1}^{N}f_{i}\cdot\delta(u_{i},k)}{\sum_{i=1}^{N}f_{i}}\Big{]}$
with $\delta(u_{i},k)=1$ if $u_{i}=k$ and zero otherwise.
S2.4 Adjusted Mutual Information (AMI)
Consider two partitions or clusterings of a set of $N$ elements, $U=\{u_{1},\ldots,u_{N}\}$ and $V=\{v_{1},\ldots,v_{N}\}$. The Adjusted Mutual Information (AMI) between $U$ and $V$ is a measure of the similarity between the two clusterings and it is corrected for the effect of agreements solely by chance [70]. It is defined as follows:
$AMI(U,V)=\frac{MI(U,V)-\mathbb{E}[MI(U,V)]}{\text{max}\{H(U),H(V)\}-\mathbb{E}[MI(U,V)]}$ (S26)
where $H(\cdot)$ and $MI(\cdot,\cdot)$ are the Shannon entropy and the mutual information respectively [71]. The AMI takes its maximum value $1$ when the two partitions are identical, and it takes its minimum value $AMI=0$ when the two partitions have MI equal to the expected value solely due to chance.
S2.5 Clustering of membership vectors
The inference procedure by bioSBM yields for each node a membership proportion vector of dimension $K$, rather than a single discrete label. To make a direct comparison with known biological annotations such as AB compartments, subcompartments [9], or the interaction profile groups (IPGs) by Spracklin et al. [15], we perform $k$-means clustering [55] of the membership proportion vectors from our model. To compare with AB compartment, we set the number of clusters to 22, while to compare with subcompartments and IPGs, we set the number of clusters to 6. Then, to evaluate the clustering results, we need to deal with the fact that the numerical labels assigned to the clusters are arbitrary, due to permutation ambiguity. To match the numerical labels found with two clustering procedures, we solve a linear assignment problem, where the goal is to maximize the overlap between clusters. We solve the matching problem using the so-called Hungarian algorithm [72].
@@
H3K27me3 (ChIP-seq) ENCSR810BDB ENCFF717ZKL
H3K4me1 (ChIP-seq) ENCSR161MXP ENCFF239FXT
H3K36me3 (ChIP-seq) ENCSR091QXP ENCFF024LGD
HCT116 H3K9me3 (ChIP-seq) ENCSR179BUC ENCFF254TIW
and H3K9ac (ChIP-seq) ENCSR093SHE ENCFF743PJP
HCT116 RAD21- H3K4me2 (ChIP-seq) ENCSR794ULT ENCFF693HVA
H4K20me1 (ChIP-seq) ENCSR474DOV ENCFF349NQI
H2AFZ (ChIP-seq) ENCSR227XNT ENCFF863EHT
H3K79me2 (ChIP-seq) ENCSR494CCN ENCFF631DLM
CTCF (ChIP-seq) ENCSR450DXU ENCFF787LAV
POLR2A (ChIP-seq) ENCSR [^0] EUU ENCFF802CGI
RAD21 (ChIP-seq) ENCSR [^0] BSB ENCFF776IXR
H3K4me3 (ChIP-seq) ENCSR668LDD ENCFF253TOF
H3K27ac (ChIP-seq) ENCSR [^0] AKP ENCFF381NDD
H3K27me3 (ChIP-seq) ENCSR [^0] AKQ ENCFF242ENK
H3K4me1 (ChIP-seq) ENCSR [^0] AKS ENCFF607SUJ
H3K36me3 (ChIP-seq) ENCSR [^0] AKR ENCFF317VHO
H3K9me3 (ChIP-seq) ENCSR [^0] APE ENCFF601JGK
K562 H3K9ac (ChIP-seq) ENCSR [^0] AKV ENCFF286WRJ
H3K4me2 (ChIP-seq) ENCSR [^0] AKT ENCFF959YJV
H4K20me1 (ChIP-seq) ENCSR [^0] AKX ENCFF605FAF
H2AFZ (ChIP-seq) ENCSR [^0] APC ENCFF621DJP
H3K79me2 (ChIP-seq) ENCSR [^0] APD ENCFF544AVW
CTCF (ChIP-seq) ENCSR [^0] AKO ENCFF979PWH
POLR2A (ChIP-seq) ENCSR031TFS ENCFF124WLE
RAD21 (ChIP-seq) ENCSR [^0] FAD ENCFF320RTQ
GM12878 in situ Hi-C 4DNES3JX38V5 4DNFI1UEG1HD
HCT116 RAD21- in situ Hi-C 4DNESJV9TH8Q 4DNFILIM6FDL
HCT116 in situ Hi-C 4DNES3QAGOZZ 4DNFIP71EWXC
K562 in situ Hi-C 4DNESI7DEJTM 4DNFITUOMFUQ
@@
Figure S1: The Evidence Lower BOund (ELBO) was used as a criterion for Bayesian model selection. For each dataset and each value for the number of communities, the variational algorithm was run multiple times with random initialization. Each run was stopped when the relative increase of the ELBO went below $10^{-5}$. The final ELBO values as a function of K was fitted with a saturating exponential of the form $a(1-e^{-K/\tilde{K}})+b$ and the turquoise band represents a saturation in the range $93.5-96.5\%$. a) shows that for the dataset we used to perform predictions (all chromosomes of the GM12878 cell line), we identify the number of communities to be between 6 and 7. b) shows that the model selection is robust to subsampling the dataset, as the model selection of the odd numbered chromosomes and even numbered chromosomes yields consistent results c) shows that the number of communities to describe the interaction structures are different in different cell lines.
Figure S2: (a) Shannon entropies of the MAP estimates of the membership vectors for all analyzed genomic regions. The peak at zero shows that a big fraction of nodes displays a single membership behavior, however, there is still a considerable degree of mixed-membershipness if one looks at the entirety of the genomic regions under study. The pattern appears to be conserved across different cell lines. (b) Blue box plots show the prediction correlations reported in Fig. 5 in main text. Red box plots represent the same predictions, but with the membership vectors reduced to the most probable (single) community assignment. (c) Examples of patterns of genomic regions with high or low entropy in membership proportions, classified with the 50th percentile of the distribution from panel (a). (d) Saddle plot computed on genome-wide Hi-C contacts in GM12878, showing contact differences between regions stratified by entropy levels.
Figure S3: Log-fold enrichment in key biochemical features for all somatic chromosomes. For each chromosome, the top heatmap represents enrichment for A/B compartments, while the bottom one is for the binary clustering obtained from bioSBM membership proportion vectors.
Figure S4: Log-fold enrichment in key biochemical features for all somatic chromosomes. For each chromosome, the left heatmap represents enrichment for subcompartments [9], the right one is for the clusters obtained with the spectral clustering algorithm from [15] using 6 groups, and the central one is for the clustering obtained from bioSBM membership proportion vectors setting the number of clusters equal to 6.
Figure S5: Hi-C maps predicted for GM12878. Bottom triangles depict experimental log O/E Hi-C maps for the GM12878 cell line. The top triangles represent model predictions. The prediction for each even-numbered chromosome is carried out by the model trained on odd-numbered chromosomes alone and vice versa.
@@
Figure S6: Correlations between model-generated maps and experimental log O/E Hi-C maps for two models with a different map from biochemical covariates to membership vectors. The turquoise plots correspond to the case where $\Gamma$ is the simple $14\times 7$ matrix representing a linear mapping, discussed in main text. The red boxes correspond to the case where $\Gamma()$ is a feed-forward deep neural network with 4 hidden layers of size $(64,64,64,32)$ with the standard ReLU (rectified linear unit) non-linear activations. While on the training set and on test chromosomes from the same cell line of the training data, the non-linear model maintains a high accuracy, for predictions on other cell lines, the correlations decrease drastically.
Figure S7: Prediction accuracies for GM12878 using a subset of covariates. The blue dots refer to Pearson correlation coefficients between experimental log O/E maps and model generated contacts using the full set of 14 covariates. The green dots refer to predictions made with H3k27ac, H3k27me3, H3k9me3, H3k79me2, H3k9ac, RAD21, POLR2A. The predictions on the even-numbered chromosomes are done by re-training the regression weights $\Gamma$ (Sec. S1), with the membership proportion vectors inferred for odd-numbered chromosomes. The reverse is done for the odd-numbered chromosomes.
Figure S8: (a) Correlations between experimental maps for HCT116 cells after RAD21-degradataion ([57]), and prediction made with or without the input RAD21 track, for each chromosome, stratified by genomic distance. (b) Correlations computed using: (i) all contacts, (ii) short-range contacts at distances shorter than 2Mbps, (iii) long-range contacts at higher distances.
Figure S9: Histograms of centromere distance for different clusters. Dotted lines represent distributions for the subcompartments defined in [9]. Solid lines correspond to the six clusters obtained using the spectral clustering algorithm from [15], matched post hoc to the most similar subcompartment (see Sec. S2.5). Dash-dotted lines indicate clusters derived by applying k-Means to the membership vectors inferred by bioSBM. The absence of clearly distinct “centromere-proximal” or “centromere-distal” clusters suggests that these groupings do not reflect trivial spatial biases, which could otherwise indicate overestimation of the number of biologically meaningful clusters.